You are the manager of a firm that produces furniture. All your clients are in Kumasi (K) and Accra (A). Suppose the monthly inverse demand function for your furniture in in Accra and Kumasi combined is 𝑃 = 1200 βˆ’ 4𝑄. Because of prohibitive cost of transporting furniture between Accra and Kumasi, you set up two plants, one in each town. The cost of producing furniture at each facility is given by 𝐢A (𝑄A ) = 8000 + 6𝑄A2 and 𝐢K(𝑄K) = 8000 + 6𝑄K2 where 𝑄 = 𝑄A + 𝑄K and 𝑄A and 𝑄K are the quantity of furniture produced at Accra and Kumasi respectively. Determine the profit maximizing amounts of electricity to produce in the two facilities, the optimal price and the profit.Β 


1
Expert's answer
2021-09-26T20:16:32-0400

"P=1200-4Q\\TR=PQ=1200Q-4Q^2"

"MR=\frac{d(TR)}{dQ}=1200-8Q\\MR=1200-8Q_A-8Q_k\\C_A=8000+6Q_A^2\\MC_A=6(2)Q_A^{(2-1)}=12Q_A\\
C_k=8000+6Q_k^2\\MC_k=6(2)Q_k^{(2-1)}=12Q_k"


"MR=MC_A\\1200-8Q_A-8Q_k=12Q_k\\1200-8Q_k=20Q_k"

Dividing both sides by 20

"60-0.4Q_k=Q_A\\MR=MC_k\\1200-8Q_A-8Q_k=12Q_k\\1200-8Q_k=20Q_k"


"Q_k=60-0.4Q_A\\Q_k=60-0.4(60-0.4Q_k)\\Q_k=60-24+0.16Q_k\\0.84Q_k=36\\Q_k=42.86\\Q_A=60-0.4(42.86)=42.86"

Optimal price"=1200-4(42.86+42.86)=857.14"

Profit"=p(Q_A+Q_k)-C_A-C_k"

"=857.14(42.86+42.86)-(8000+6(44.86^2))\\=35428.57"